# Can Large Reasoning Models do Analogical Reasoning under Perceptual Uncertainty?

### v1

@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 (OpenAI et al., 2024), Gemini (Gemini Team et al., 2024), and Claude (Anthropic, 2024) have demonstrated great proficiency in generating fluent and contextually relevant text. However, their capabilities in more complex domains such as reasoning and planning have been shown to be brittle even in simple tasks (Gendron et al., 2024; Wu et al., 2024), fail to attain levels of general abstract reasoning comparable to humans (Odouard and Mitchell, 2022; Thomm et al., 2024; Lewis and Mitchell, 2025; Camposampiero et al., 2023) and may be in some cases a result of data contamination (Roberts et al., 2023; Mirzadeh et al., 2025). To mitigate this issue, the attention has shifted from training-time to inference-time compute scaling. This resulted in the development of a new generation of systems, dubbed Large Reasoning Models (LRMs), that can dynamically allocate variable compute time during inference based on the input query. Unlike LLMs, which behaved mostly like approximate retrievers, LRMs such as OpenAI o1 (OpenAI, 2024), OpenAI o3, DeepSeek R1 (DeepSeek-AI et al., 2025), and Qwen QwQ (Team, 2024) approach reasoning tasks by exploring the space of solutions through pseudo-actions using Chain-of-Thought (CoT) (Wei et al., 2022) tokens.
@@
Figure 1: Analogical reasoning under perceptual uncertainty. In this figure, we highlight all the different axes of generalization and robustness to uncertainty, which I-RAVEN-X stresses. Compared to standard I-RAVEN (a), I-RAVEN-X (b) involves more panels per row (10 vs. 3) and larger attribute dynamic ranges (up to 100×100$\times$100 × more values per attribute). This work introduces uncertainty in the reasoning process (c) through confounders (such as panels’ background and color patterns within objects) and smoothening the attribute values’ distributions (displayed on the right for the panel in position $(1,10)$). The RPM rules used in this example are constant for shape, distribute right for color, constant for size, and progression for number. We adopt a visual representation of the panels and their attributes for clarity of explanation; in practice, however, our dataset is purely symbolic and has not been extended yet to the visual domain.
While LRMs achieved remarkable performance on many reasoning benchmarks in their preferred textual domain, they still have not been successful in other modalities, such as vision (Mitchell et al., 2024; Jiang et al., 2024; Cao et al., 2024; Ahrabian et al., 2024; Zhang et al., 2024b). Hence, it has become a standard practice to assume the availability of an oracle perception and prompt LRMs with ideal and discrete symbolic transcriptions of the test examples (Webb et al., 2023; Hu et al., 2023; Hersche et al., 2025). However, the assumption of an oracle visual perception bypasses crucial steps in visual abstract reasoning. First, the oracle perception takes the knowledge about the set of attributes needed to perform the reasoning task for granted. As a result, background variables that do not influence the answer prediction are automatically filtered out, effectively eliminating a crucial step in the reasoning process. Second, the oracle perception usually provides the attribute values with full confidence, i.e., the distribution over the attributes is a degenerate probability mass function (PMF). This assumption is highly unrealistic since any neural perception will always feature some degree of uncertainty in its outputs. In summary, a general reasoning method should not only be able to perform reasoning on perfect input representations but also handle uncertainties with respect to the number of variates and their distribution.
This paper proposes a more thorough method to benchmark LRMs on abstract visual reasoning (see Figure 1). Specifically, we focus on solving Raven’s progressive matrices (RPMs, Raven et al. (1938)), a visual abstract analogical reasoning task that is commonly used to test the fluid intelligence of humans and, more recently, of machines too (Barrett et al., 2018; Zhang et al., 2019; Hu et al., 2021; Jiang et al., 2024). To avoid potential data leakage from the model’s pre- and post-training stages (Mirzadeh et al., 2025), we propose a fully symbolic, generative methodology to evaluate LRMs in a realistic scenario. The resulting dataset, dubbed I-RAVEN-X, evaluates the reasoning capabilities from the following aspects:
  • Productivity: larger context matrix sizes are introduced, e.g., 3$\times$10 instead of 3$\times$3;
  • Systematicity: larger dynamic ranges for the attribute values are introduced, e.g., 1000 attribute values instead of 10;
  • Robustness to confounding factors: the set of generative attributes in RPM is augmented with randomly sampled values, which do not contribute to the reasoning;
  • Robustness to non-degenerate value distributions: the distributions of the input values corresponding to the generative factors are smoothened.
While an initial attempt evaluated the productivity using I-RAVEN-X (Hersche et al., 2025), this work proposes a more elaborate benchmarking using LRMs with an additional focus on robustness. We leverage I-RAVEN-X to perform an exhaustive analysis on two state-of-the-art (SOTA) LRMs, OpenAI’s o3-mini and DeepSeek R1. Compared to I-RAVEN (Hu et al., 2021), I-RAVEN-X’s increased matrix size and dynamic range reduce o3-mini’s task accuracy from $86.6\%$ to $81.0\%$. Moreover, the task accuracy further degrades when additionally introducing 10 confounding factors ($69.8\%$) and the smoothened value distribution ($75.6\%$). Finally, when I-RAVEN-X combines all its productivity and robustness measures, the accuracy sharply decreases to $17.0\%$. Similar trends are observed for R1, experiencing an overall drop from $80.6\%$ to $23.2\%$.
As a potential remedy to LRM’s susceptibility to perceptual uncertainty, we show that neuro-symbolic probabilistic abductive reasoning methods can naturally support the uncertainty from smoothened distributions. In addition, we present a novel entropy-based confidence metric that allows probabilistic abductive reasoning methods to filter confounders in the decision process. We evaluate the new confidence metric on one of the SOTA approaches of this family (ARLC, Camposampiero et al. (2024)). In the most difficult I-RAVEN-X setting (mix of confounders and smoothening noise), ARLC achieves significantly stronger accuracy ($88.3\%$) compared to LRMs. Moreover, ARLC can retain high accuracy even in very harsh signal-to-noise ratios (SNRs) conditions (up to $-20$ dB).
2 Background and Related Works
Analogical Reasoning Benchmarks
@@ 2 Background and Related Works
A wide range of benchmarks to assess human-like fluid intelligence and abstract reasoning has been proposed in the past decade (Bilker et al., 2012; Cherian et al., 2023; Chollet, 2019; Niedermayr et al., 2024). Raven’s Progressive Matrices (RPM) (Raven et al., 1938; Carpenter et al., 1990; Bilker et al., 2012) is one of the most prominent among them due to its extensive to benchmark for abstract reasoning, analogy-making, and out-of-distribution (OOD) testing (Benny et al., 2021; Hu et al., 2021; Małkiński and Mańdziuk, 2025; Mitchell, 2021; Zhang et al., 2019). RAVEN (Zhang et al., 2019) represented one of the first attempts to build a dataset in the context of RPM that aimed at associating vision with structural, relational, and analogical reasoning in a hierarchical representation. I-RAVEN (Hu et al., 2021) (Figure 1a) improved RAVEN, proposing a new generation algorithm based on attribute bisection trees. This ensured that candidate panels are sampled from an unbiased candidate set, avoiding shortcut solutions that were possible in the original dataset. I-RAVEN-X (Hersche et al., 2025) extended I-RAVEN, introducing a parameterizable number of columns and a dynamic range of attribute values that allow testing the generalization of analogical reasoning to longer reasoning chains and an increased number of concepts. In Figure 1b, this can be identified by the larger number of columns and a wider range of attributes (such as the color and the number of objects), respectively. In addition, the dataset was narrowed down to a single constellation (center, containing only one object per panel) which was observed to be at the same time a strong test for a wide range of logical and arithmetic skills and unexpectedly challenging for LLMs (Hersche et al., 2025).
Large Reasoning Models
@@
Recent research has focused on training LLMs to exhibit human-like reasoning (OpenAI, 2024), yet a major obstacle remains the scarcity of annotated, step-by-step reasoning data. To overcome this issue, researchers started transitioning from expensive human annotations to LLM-driven search algorithms that automatically generate accurate reasoning trajectories through external verification (Luo et al., 2024) and RL-based techniques (Zhang et al., 2024a; Shao et al., 2024). Moreover, scaling test-time computation was also shown to be useful to refine intermediate reasoning steps, thereby further improving accuracy on reasoning tasks (Snell et al., 2024). Together, the combination of RL-driven train-time scaling and search-based test-time scaling paved the way for a new generation of systems, named Large Reasoning Models (LRMs), with significantly enhanced reasoning performance Xu et al. (2025). However, contrary to LLMs (Webb et al., 2023; Hu et al., 2023; Hersche et al., 2025; Moskvichev et al., 2023; Mitchell et al., 2024; Lewis and Mitchell, 2025), the analogical reasoning abilities of LRMs have not yet been extensively evaluated, with only Latif et al. (2024) showing limited results on a subset of RAVEN.
Neuro-Symbolic Architectures for RPM
@@
Improving on monolithic deep learning models (Wu et al., 2020; Benny et al., 2021), neuro-symbolic architectures implementing abductive reasoning (Magnani, 2009) achieved remarkable success, scoring SOTA results on this analogical reasoning test. Initially introduced with the PrAE learner (Zhang et al., 2021), this approach was further improved by the NVSA model (Hersche et al., 2023). In NVSA, the probabilistic reasoning was implemented via distributed representations and operators of vector-symbolic architectures (VSAs) (Gayler, 2003; Kanerva, 2009; Plate, 1995). VSAs, besides their computational and scalability benefits, provide a common language with neural networks for better interface and deeper integration. Both PrAE and NVSA are examples of Neuro | Symbolic (type 3) according to Kautz (2022); that is, they are systems composed of a neural vision module that interacts with a static symbolic reasoning system through a well-defined interface. Follow-up works extended these systems, mostly moving from pure knowledge representations to more trainable architectures that can learn from examples to reason and improve their expressiveness  (Zhang et al., 2022; Camposampiero et al., 2024; Sun et al., 2025). Some of them, as ARLC (Camposampiero et al., 2024), could be classified as Neuro[Symbolic] systems (type 6) since the reasoning rules are learned fully differentiably from a generic rule template encoded in distributed representations, and it is capable of combinatorial reasoning by exploiting computation-in-superposition.
@@ 3 Integrating perceptual uncertainty into I-RAVEN-X
Contrary to the standard I-RAVEN, I-RAVEN-X is a fully symbolic benchmark that evaluates abstract reasoning under the assumption of an oracle perception. This assumption stems from the observation that using the original visual inputs to prompt multi-modal LLMs performed significantly worse compared to noiseless symbolic transcriptions of the test examples (Mitchell et al., 2024; Jiang et al., 2024; Cao et al., 2024; Ahrabian et al., 2024; Zhang et al., 2024b). However, it represents a rather strong assumption since it neglects the uncertainty that would necessarily result from the extraction of those attributes in real-world scenarios and its influence on the analogical reasoning process. In this work, we propose an extension of I-RAVEN-X to overcome it, augmenting the original dataset by:
@@ 3.1 Confounding attributes
Confounding attributes represent properties and patterns that can be extracted from the visual inputs by a front-end perception module but are not relevant to the reasoning process. This could be the case, for instance, when the attributes are extracted by unsupervised vision models such as Variational Autoencoders (Kingma and Welling, 2022) or even a multi-modal LLM that is prompted to extract the attributes. In Figure 1c, confounding attributes are represented by the background of the input panels and the color patterns, which sometimes appear inside the objects. While the original RAVEN dataset includes noise attributes (e.g., the orientation), we argue that these are not real confounders and that their naming is misleading, as they do not add any noise to the RPM tests (we argument this more extensively in Appendix C). In I-RAVEN-X, this is practically implemented by extending the set of original attributes of each panel with an arbitrary number of confounding factors uniformly sampled in the interval $[0,m-1]$, where $m$ is the dynamic range of the attributes in the experiment. For large enough $m$, the probability of sampling values that fit a valid rule is negligible, and hence, confounders do not introduce ambiguities in the choice of the answer panel. However, they linearly reduce the SNR in the reasoning process and require models to filter the noisy attributes.
@@ 3.2 Smooth attribute values’ distributions
We deviate from the original I-RAVEN-X degenerate attributes’ distributions and introduce variance, which allows us to test the robustness of the models when reasoning with uncertain attributes’ values. Figure 1 highlights this relaxation, from one-hot PMFs of the standard I-RAVEN-X (Figure 1b) to the distributed PMFs of our proposed extension (Figure 1c). In practice, we smoothen the original attributes’ distributions using either a Gaussian filter or with a three-bins strategy, where the probability of the true value $T$ is $p(T)\sim\mathcal{U}(p_{L},1),p_{L}>0.5$ and the probabilities of its two neighboring values are $p(N_{1})\sim\mathcal{U}(0,1-p(T))$ and $p(N_{2})=1-p(T)-p(N_{1})$. Note that the motivation behind the three-bins strategy is to introduce variance with minimal additional for LRMs prompt complexity.
@@ 4 Solving RPMs with LRMs and NeSy Probabilistic Abductive Models
4.1 Large Reasoning Models (LRMs)
[…]
We adopt the same evaluation framework used in Hersche et al. (2025) to benchmark LRMs. Contrary to their analysis, however, we focus our investigation on entangled prompts (providing all the attributes’ values in a single prompt rather than having a single prompt per attribute). We are forced to choose this setting because, despite performing worse compared to using disentangled prompts (where one separate prompt is used for each attribute), successive experiments on confounders would have otherwise become trivial. Furthermore, we move from a predictive approach (where the model has to generate the missing panel) to a discriminative approach (where the model is given a list of candidates and required to choose one among them) (Gendron et al., 2024; Hersche et al., 2025). This choice stems from observing, in the early stages of our evaluation, that LRMs can sometimes pick up valid relations in the input matrices (for instance, relations between the binary encodings of values) which are however not part of the set of rules used in RPM. Unlike the generative approach, the discriminative approach implicitly biases the model into evaluating only the rules defined in RPM without explicitly revealing them, hence reducing the aforementioned issue. For more details on the task and prompts, refer to Appendix A.
Self-consistency (Wang et al., 2023; Lewkowycz et al., 2022) and attributes’ scaling (Hu et al., 2023) were also dropped in the experiments with LRMs. Moreover, no in-context examples of the tasks (Brown et al., 2020) were provided since they were previously observed to be hurtful for LRMs (DeepSeek-AI et al., 2025). We also restrict the investigation to a subset of 500 randomly sampled RPM tests in both I-RAVEN and I-RAVEN-X (due to budget constraints), which we observed to be representative enough of the entire test set.

### v2

@@ 1 Introduction
Figure 1: Analogical reasoning under perceptual uncertainty. This figure highlights all the different axes of generalization and robustness to uncertainty, which I-RAVEN-X stresses. Compared to standard I-RAVEN (a), I-RAVEN-X (b) involves more panels per row (10 vs. 3) and larger attribute dynamic ranges (up to 100×100$\times$100 × more values per attribute). This work (c) introduces uncertainty in the reasoning process through confounders (such as panels’ background and color patterns within objects) and smoothening the attribute values’ distributions (displayed on the right for the panel in position $(1,10)$). We adopt a visual representation of the panels and their attributes for clarity of explanation; in practice, however, our dataset is purely symbolic and has not been extended yet to the visual domain.
Large Language Models (LLMs) such as GPT-4 (OpenAI et al., 2024), Gemini (Gemini Team et al., 2024), and Claude (Anthropic, 2024) have demonstrated great proficiency in generating fluent and contextually relevant text. However, their capabilities in more complex domains, such as reasoning and planning, have been shown to be brittle even in simple tasks (Gendron et al., 2024; Wu et al., 2024), fail to attain levels of general abstract reasoning comparable to humans (Odouard and Mitchell, 2022; Thomm et al., 2024; Lewis and Mitchell, 2025; Camposampiero et al., 2023) and may be in some cases a result of data contamination (Roberts et al., 2023; Mirzadeh et al., 2025). To mitigate this issue, the attention has shifted from training-time to test-time compute scaling. This resulted in the development of a new generation of systems, dubbed Large Reasoning Models (LRMs), that can dynamically allocate variable compute at test time based on the input query. Unlike LLMs, which behaved mostly like approximate retrievers, LRMs such as OpenAI’s o1 and o3 (OpenAI, 2024), DeepSeek R1 (DeepSeek-AI et al., 2025), and Qwen QwQ (Qwen Team, 2024) approach reasoning tasks by exploring the space of solutions through pseudo-actions using Chain-of-Thought (CoT) (Wei et al., 2022) tokens. Contrary to LLMs (Webb et al., 2023; Hu et al., 2023; Hersche et al., 2025; Moskvichev et al., 2023; Mitchell et al., 2024; Lewis and Mitchell, 2025), however, the analogical reasoning abilities of LRMs have not yet been extensively evaluated (l1), with only Latif et al. (2024) showing limited results on it. Furthermore, we identify a critical limitation of recent works that benchmarked language models on visual analogical reasoning. Due to their poor performance in the visual domain (Mitchell et al., 2024; Jiang et al., 2024b; Cao et al., 2024; Ahrabian et al., 2024b; Zhang et al., 2024b), assuming the availability of an oracle perception and prompt LLMs with discrete symbolic transcriptions of the test examples has become a standard practice (Webb et al., 2023; Hu et al., 2023; Hersche et al., 2025). However, this assumption bypasses crucial steps in visual analogical reasoning. First, the oracle perception takes the knowledge about the set of attributes needed for the reasoning task for granted, avoiding the step of filtering attributes that are irrelevant to the answer prediction (l2). Second, the oracle perception usually provides the attribute values with full confidence, an unrealistic assumption since any perception model will always feature some degree of uncertainty (l3). In order to tackle these issues, this paper makes the following contributions:
  1. 1.
    In Section 3, we introduce a more thorough method to benchmark analogical reasoning in LRMs, extending the fully-symbolic dataset I-RAVEN-X (Hersche et al., 2025) with simulated perceptual uncertainty given by confounding attributes (tackling l2) and uncertainty on the reasoning attributes’ values (tackling l3), as shown in Figure 1. The resulting dataset, I-RAVEN-X, improves on the standard I-RAVEN (Hu et al., 2021) by allowing to test a) productivity (longer reasoning relations), systematicity (more values for each attribute), robustness to confounding factors (the set of generative attributes in RPM is augmented with randomly sampled values, which do not contribute to the reasoning), and robustness to non-degenerate value distributions (uncertainty associated to the value of each attribute).
  2. 2.
    In Section 5, we present the first exhaustive analysis on two state-of-the-art (SOTA) LRMs, OpenAI’s o3-mini and DeepSeek R1, on symbolic analogical reasoning using the newly-introduced I-RAVEN-X (addressing l1). Despite showing significant improvement over LLMs, LRMs are affected by a a sharp decay in accuracy when simultaneously evaluated on all generalization dimensions (productivity, systematicity, and robustness), with o3-mini and R1 dropping from $86.6\%$ to $17.0\%$ and $80.6\%$ to $23.2\%$, respectively, substantially approaching random chance.
  3. 3.
    In Section 4.2, we propose a novel entropy-based regularizer integrable in abductive reasoning frameworks to improve their robustness to noisy features, implicitly allowing them to filter confounders in the decision process. We implement this regularizer in ARLC (Camposampiero et al., 2024) and show that it helps retaining high accuracy in harsh signal-to-noise ratio (SNR) conditions (up to $-20$ dB). Furthermore, we show that ARLC, as a neuro-symbolic probabilistic abductive reasoner supporting reasoning under uncertainty, achieves significantly stronger accuracy ($88.3\%$) compared to LRMs.
2 Related works
Analogical reasoning benchmarks.
@@ 2 Related works
A wide range of benchmarks to assess abstract reasoning has been proposed in the past decade (Bilker et al., 2012; Cherian et al., 2023; Chollet, 2019; Niedermayr et al., 2024). Raven’s Progressive Matrices (RPM) (Raven et al., 1938; Carpenter et al., 1990; Bilker et al., 2012) is one of the most prominent among them due to its extensive use to benchmark for abstract reasoning, analogy-making, and out-of-distribution (OOD) testing (Benny et al., 2021; Hu et al., 2021; Małkiński and Mańdziuk, 2025; Mitchell, 2021; Zhang et al., 2019). RAVEN (Zhang et al., 2019) represented one of the first attempts to build a dataset in the context of RPM that aimed at associating vision with structural, relational, and analogical reasoning in a hierarchical representation. I-RAVEN (Hu et al., 2021) (Figure 1a) improved RAVEN, proposing a new generation algorithm based on attribute bisection trees. This ensured that candidate panels are sampled from an unbiased candidate set, avoiding shortcut solutions that were possible in the original dataset. I-RAVEN-X (Hersche et al., 2025) extended I-RAVEN, introducing a parameterizable number of columns and a dynamic range of attributes that allow testing the generalization of analogical reasoning to longer reasoning chains and an increased number of concepts. In Figure 1b, this can be identified by the larger number of columns and a wider range of attributes (such as the color and the number of objects), respectively. In addition, the dataset was narrowed down to a single constellation (center, containing one object per panel) which was observed to be simultaneously a strong test for a wide range of logical and arithmetic skills and unexpectedly challenging for LLMs (Hersche et al., 2025).
Large reasoning models.
@@
Recent research has focused on training LLMs to exhibit human-like reasoning (OpenAI, 2024), yet a major obstacle remains the scarcity of annotated, step-by-step reasoning data. To overcome this issue, researchers started transitioning to LLM-driven search algorithms that automatically generate accurate reasoning trajectories through external verification (Luo et al., 2024) and RL-based techniques (Zhang et al., 2024a; Shao et al., 2024). Moreover, scaling test-time computation was also shown to be useful to refine intermediate reasoning steps and further improve accuracy on reasoning tasks (Snell et al., 2024). Together, the combination of RL-driven train-time scaling and search-based test-time scaling paved the way for a new generation of systems, Large Reasoning Models (LRMs), with significantly enhanced reasoning performance (Xu et al., 2025).
Neuro-symbolic architectures for RPM.
@@
Improving on monolithic deep learning models (Wu et al., 2020; Benny et al., 2021), neuro-symbolic architectures implementing abductive reasoning (Magnani, 2009) achieved remarkable success, scoring SOTA results on this analogical reasoning test. Initially introduced with the PrAE learner (Zhang et al., 2021), this approach was further improved by the NVSA model (Hersche et al., 2023). In NVSA, the probabilistic reasoning was implemented via distributed representations and operators of vector-symbolic architectures (VSAs) (Gayler, 2003; Kanerva, 2009; Plate, 1995). VSAs, besides their computational and scalability benefits, provide a common language with neural networks for better interface and deeper integration. Both PrAE and NVSA are examples of Neuro | Symbolic networks (Kautz, 2022); that is, they are systems composed of a neural vision module that interacts with a static symbolic reasoning system through a well-defined interface. Follow-up works extended these systems, mostly moving from pure knowledge representations to more trainable architectures that can learn from examples to reason and improve their expressiveness  (Zhang et al., 2022; Camposampiero et al., 2024; Sun et al., 2025). Some of them, as ARLC (Camposampiero et al., 2024), could be classified as Neuro[Symbolic] systems (Kautz, 2022) since the reasoning rules are learned in a differentiable fashion from a generic rule template encoded in distributed representations, and it is capable of combinatorial reasoning by exploiting computation-in-superposition.
@@ 3 Integrating perceptual uncertainty into I-RAVEN-X
Contrary to the standard I-RAVEN, I-RAVEN-X is a fully symbolic benchmark that evaluates analogical reasoning under the assumption of an oracle perception. However, this is a rather strong assumption that neglects the uncertainty that would necessarily result from the extraction of those attributes in real-world scenarios and their influence on the analogical reasoning process. In this work, we propose an extension of I-RAVEN-X to overcome it, augmenting the original dataset by:
@@ 3.1 Confounding attributes
Confounding attributes represent properties and patterns that can be extracted from the visual inputs by a front-end perception module but are not relevant to the reasoning process. This could be the case, for instance, when the attributes are extracted by unsupervised vision models such as Variational Autoencoders (Kingma and Welling, 2013) or even a multi-modal LLM that is prompted to extract the attributes. In Figure 1c, confounding attributes are represented by the background of the input panels and the color patterns, which sometimes appear inside the objects. While the original RAVEN dataset includes noise attributes (e.g., the orientation), we argue that these are not real confounders as they do not add any noise to the RPM tests (refer to Appendix C for a more extensive discussion). In I-RAVEN-X, we integrate confounders by extending the set of original attributes of each panel with an arbitrary number of additional attributes uniformly sampled in the interval $[0,m-1]$, where $m$ is the range of the attributes’ values. For large enough $m$, the probability of sampling values that fit a valid rule is negligible, and hence, confounders do not introduce ambiguities in the choice of the answer panel. However, they linearly reduce the SNR in the reasoning process, requiring models to implement strategies to filter out noisy input components.
@@ 3.2 Smooth attribute values’ distributions
We deviate from the original I-RAVEN-X degenerate attributes’ distributions and introduce variance, which allows us to test the robustness of the models when reasoning with uncertain attributes’ values. Figure 1 highlights this relaxation, from one-hot probability mass functions (PMFs) of the standard I-RAVEN-X (Figure 1b) to the distributed PMFs of our proposed extension (Figure 1c). In practice, we smoothen the original attributes’ distributions using either a Gaussian filter or with a three-bins strategy, where the probability of the true value $T$ is $p(T)\sim\mathcal{U}(p_{L},1),p_{L}>0.5$ and the probabilities of its two neighboring values are $p(N_{1})\sim\mathcal{U}(0,1-p(T))$ and $p(N_{2})=1-p(T)-p(N_{1})$. Note that the motivation behind the three-bins strategy is to introduce variance with minimal additional cost for LRMs’ prompt complexity.
@@ 4 Solving RPMs with LRMs and NeSy probabilistic abductive models
4.1 Large reasoning models
We focus our study on the two most prominent SOTA LRMs available to date: the closed-source OpenAI o3-mini model and the open-source DeepSeek R1 model (DeepSeek-AI et al., 2025) (together with its distilled version based on Llama 70B). In Appendix B, we include an additional (limited) comparison between OpenAI o3-mini and its predecessor, OpenAI o1. However, since the performance of the o3-mini model was on par with o1 while costing only a fraction (14$\approx$ 14$\times$ 14 × less), we decided to experiment only with o3-mini.
We adopt the same evaluation framework used in Hersche et al. (2025) to benchmark LRMs. Contrary to their analysis, however, we focus our investigation on entangled prompts (providing all the attributes’ values in a single prompt rather than having a single prompt per attribute). We are forced to choose this setting because, despite performing worse compared to using disentangled prompts (where one separate prompt is used for each attribute), successive experiments on confounders would have otherwise become trivial. Furthermore, we move from a predictive approach (where the model has to generate the missing panel) to a discriminative approach (where the model is given a list of candidates and required to choose one among them) (Gendron et al., 2024; Hersche et al., 2025). This choice stems from observing, in the early stages of our evaluation, that LRMs can sometimes pick up valid relations in the input matrices (for instance, relations between the binary encodings of values) which are however not part of the set of rules used in RPM. Unlike the generative approach, the discriminative approach implicitly biases the model into evaluating only the rules defined in RPM without explicitly revealing them, hence reducing the aforementioned issue. For more details on the task, models, and prompting strategy, refer to Appendix A.